OpenStudy (anonymous):

If * is a binary operation defined by a*b = minimum of {a,b} or a min b, find the value of 10*8

OpenStudy (anonymous):

what is smaller, 10 or 8?

OpenStudy (anonymous):

\(10*8=\min\{10,8\}=?\) which is the minimum? ;-)

OpenStudy (anonymous):

\[a*b=\min(a,b)\] \[10*8=\min(10,8)\] What's the smaller number?

OpenStudy (anonymous):

Haha the answers to this question.

OpenStudy (anonymous):

Hehe the responses to this query.

OpenStudy (anonymous):

Erm. 8

OpenStudy (anonymous):

With questions like this don't just panic as soon as you see the word 'binary' or operation' , just read the question and understand it, 9 times out of 10 it's easy as..

OpenStudy (anonymous):

multiplication?

OpenStudy (anonymous):

multiplication is one example of a binary operation , but not the only one "binary" because it takes "2" numbers to multiply, and combines them in to one number \[3\times 4=12\] addition is also "binary" \[3+4=7\] two number become one

OpenStudy (anonymous):

in this case the "binary" operation is not one you are used to it says "take two numbers, compare them, pick the smallest"

OpenStudy (anonymous):

so it is not multiplication or addition, but something else. if we call the operation \(\circ\) then for example \(3\circ 5=3\) because \(3\) is the smaller of the two

OpenStudy (anonymous):

Ohh..

OpenStudy (anonymous):

So the answer is 8?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the answer to the question is very easy the hard part is understanding what it is is asking that is why @sarahusher wrote her answer above, understanding the question is the key to solving it

OpenStudy (anonymous):

Ohh....okay! Thank chuu!